An airline knows that 6 % of its booking won't show up at the airport to travel.

If in a flight of 250 passengers, 260 people make a booking:

a) What's the probabality that all who show up will eventually travel

b) What's the probability that exactly 240 people will travel

c) How many should the bookings be for the airline to know with a probability of 99% that the flight will be full
Answer #1

P(X) = nCx p*(1-p)*-*

x = no. of people who show up

n = 260

p = P(show up) = 1 - 0.06 = 0.94

1-p = 0.06

a.

probabality that all who show up will eventually travel

= P(x<=250)

= 0.9524

b.

probability that exactly 240 people will travel

P(240)

= 260C240 * 0.94^240 * 0.06^20

= 0.0503

c.

P(full) = P(x>=250)

on trial and error we find for n = 276

we get P(x>=250) = 0.99

(please UPVOTE)